Commercial brass, an alloy of Zn and Cu, reacts with hydrochloric acid to form zinc chloride and hydrogen gas. Copper does not react with HCl. When 0.6503 g of a brass sample is reacted with excess HCl, 0.1087 g ZnCl2 is eventually isolated. What is the percent (%) composition of brass by mass?


1
Expert's answer
2022-03-08T13:42:02-0500

Zn(s)"+" 2HCl(aq)"\to" ZnCl2(aq)"+" H2(g)


Moles of ZnCl2 "=\frac{0.1087}{136.286}"


"=7.976\u00d710^{-4}" Moles


Mass of Zn "=7.976\u00d710^{-4}\u00d765.38"

"=0.05215g"



"\%" of Zn "=\frac{0.05215}{0.6503}=8.02\%"


"\%" of Cu "=100-8.02"

"=91.98\%"